OpenStudy (anonymous):

If g(x) = 3x + 5, what is g^–1(x)?

OpenStudy (anonymous):

x = 3y + 5 (x - 5)/3 = y

OpenStudy (akshay_budhkar):

y=3x+5 y-5=3x x=(y-5)/3 So inverse is y=(x-5)/3

OpenStudy (anonymous):

function says 1) multiply by 3 2) add 5 inverse will say 1) subtract 5 2) divide by 3 just like if you were going to solve \[3x+5=2\]

OpenStudy (anonymous):

so you get \[f^{-1}(x)=\frac{x-5}{3}\]